how do you calculate pH of super concentrated acid (such as 98% sulphuric acid or 99,5% ethanol)?
Answer #1

Ans: The pH can be calculated for the sulfuric acid only, as the ethanol is neutral organic alcohol and pH is not applicable to it.

Now, to calculate the pH of the super concentrated sulfuric acid, first we should know the moles of sulfuric acid present in one litre volume. For that we should first calculate the Mass (M) of sulfuric acid in 1 L volume based upon the % purity and its density.

Density of the super concentrated H2SO4 = 1.84 g/cm3 (from Litt.)

Purity of the super concentrated acid = 98 %

Molecular weight (molar mass) of the sulfuric acid = 98.1 g/mol

Mass of acid in 1L volume = 1.84 * 1000 * 0.98 = 1803.2 g

Number of moles in the above sulfuric acid weight

= weight of sulfuric acid / molar mass of the acid = 1803.2 / 98.1 = 18.38 Moles

Hence, the concentration of [H+] = 18.38 x 2 = 36.76 mol/L (sulfuric acid molecule has two protons)

pH = -log [H+]

pH = -log [36.76]

pH = 1.57